OpenStudy (anonymous):

What is the change that occurs to the parent function f(x) = x2 given the function f(x) = -2x2 + 2.

OpenStudy (anonymous):

@mariah23 Do you mean f(x) = -2x^2 + 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mariah23 Okay so you need to find the values of a, h, and k. Do you know what they are?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@mariah23 a, h, and k will tell you how the change occurs. a will dictate how wide or thin your function is h will tell you how many units it moves left or right on the x axis. k will dictate how many units you function moves up and down the y axis on your graph. Usually the function that is changing you parent function would be set up like f(x) = a(x-h)^2 + k Have you seen this before?

OpenStudy (anonymous):

@mariah23 When this f(x) = a(x-h)^2 + k is simplified you get ax^2-hx+k. Since your function is f(x) = - 2x^2 + 2, a = -2 and k = +2 There is no h in this equation so it is = to 0. ANSWER: Since A is less than 0 it will be reflected over the x axis and stretched by a factor of 2. K will move up 2 units on the y axis. does this help?